Hi can someone please help me solve for this problem, it's from the Financial Management: Principles and Application book by Keown ch. 15, study question 15-13A:
(Break-even point and operating leverage) Allison Radios manufactures a complete line of radio and communication equipment for law enforcement agencies. The average selling price of its finished product is $180 per unit. The variable cost for these same units is $126. Allison Radios incurs fixed costs of $540,000 per year.
a. What is the break-even point in units for the company?
b. What is the dollar sales volume the firm must achieve in order to reach the break-even point?
c. What would be the firm’s profit or loss at the following units of production sold:
12,000 units? 15,000 units? 20,000 units?
d. Find the degree of operating leverage for the production and sales levels given in part (c).
Thank you

Answer by rfer(10417)

A) 180x=540000+126x
180x-126x=540000
54x=540000
x=10000
----------------------
B) 10000*126+540000=1,800,000
-----------------------
C)12000*180=2160000-540000-1512000=108000
15000*180=2700000-540000-1890000=270000
20000*180=3600000-540000-2520000=540000
-------------------------
D) I will let you do this part.